OpenStudy (anonymous):

a rectangular aquarium is 2/3 filled with water and then is tilted on its side until the water level coincides with one edge of the bottom. During the tilting process, 6 quarts of water are poured out. What is the volume of the tank in quarts? (a) 12 (b) 24 (c) 36 (d) 48 (e) 64

OpenStudy (anonymous):

Volume of aq = 36. when tilted as described, half of tank is full of water. Initially it was 2/3. So what poured out was 2/3-1/2 = 1/6th of the tank. And that is 6 quarts. So total must be six times what poured out.

OpenStudy (shane_b):

The tank is 2/3rds full and when tilted to where the level coincides with one bottom edge, the leftover water takes up 1/2 the volume. \[\frac{2}{3} V - \frac{1}{2}V = 6\]Now just solve for V

OpenStudy (anonymous):

v =36

OpenStudy (anonymous):

thanks guys